OpenStudy (anonymous):

Solve the fractional equation: 1/(x+2) - 1/(x-2) = 4/3

OpenStudy (anonymous):

((x-2)-(x+2))/(x^2-4)=4/3 -4/(x^2-4)=4/3 cancelling 4 on both sides and taking reciprocal on both sides -(x^2-4)=3 -x^2+4=3 -x^2=-1 x^2=1 x=1 or x=-1